Child pages
  • Pt 12.12.2020

Measuring Pt-on-CdS and rGO NPs, Pt L3 edge.

Sample 2 shows a very small edge jump in fluorescence and nothing in transmission.

Sample1  looks better.

We don't use filters. Pt was calibrated to foil for Akhil's measurements.

Beam size : slits wide open (hor 3 * ver 4)

Flux 2E+12 photons

3und harmonic

Mono?


Sample2: very weak signal. I put all foils into a stack.

300 s scans, 200 to 1000 eV. Scans 1 to 6 are at vertical position -19.2. There was a weird bump, different from sample1; i moved to -18.7 to cross check.

The bump is still there for scan 7.

I will stay at this point for the night.


Later notes:

For sample 1 it is quite straightforward: probably all measurements were done in the same spot (judging by the fact that non-normalised spectra all overlap closely in preedge and postedge regions).

Sample1_0001 is the first 1000 eV scan; sample1_0002 is an attempt at bigger E range (1200), but there i saw that there's an edge of another element coming right after 1000 eV (Se i think), so sample1_0003 is stopped. Starting from sample1_0004 there are just repetitions with the same parameters as sample1_0001. There are 28 scans in total. Plotting them together shows that there are some changes. Edge position stays where it it. But right shoulder of the white line gets broader an the line itself grows in intensity. In R-space there is one shell (Pt-Pt) i guess, and its intensity reduces in the course of the measurement session. So there's clearly radiation damage, but it's quite slow.

Concerning sample2.

Quantity of Pt in sample2 was smaller. First, we found a foil with darkest spot and tried to measure there. The result is sample2_test_0001. The edge jump was very small. So we tried to make a stack of all foil pieces. These measurements have _ml in the filename. First, after putting the new sample in, i forgot to return the detector to the right spot, so scans ...ml_0001 to _0005 are not usable: there edge jump is almost missing there. Starting from ...ml_0006 the detector is in place. After getting a spectrum ...ml_0006 at vertical position -19.2, i moved to -18.7. All scans starting from ...ml_0007 are done in the same spot. I left the measurements for the night, so there are 126 scans in total.

Comparing scans at the beginning and at the end, the radiation damage direction looks the same as for sample1: there is a shift of the right shoulder of the white line to higher energy and growth of white line intensity.

Comparing sample1_0001 and sample2_test_0001 one can see that edge position looks the same. However, for sample2 the right shoulder of the white line is already shifted a bit to higher energies compared to fresh sample1. However, it has to be noted that low signal from the sample2 didn't allow us to use reduced flux for searching for sample position. By the time sample2_test_0001 was measured, it was already illuminated for some time.

  • No labels